A particle with mass 1.09 kg oscillates horizontally at the end of a horizontal spring. A student measures an amplitude of 0.891 m and a duration of 131 s for 68 cycles of oscillation. Find the frequency, f, the speed at the equilibrium position, vmax, the spring constant, k, the potential energy at an endpoint, Umax, the potential energy when the particle is located 50.3% of the amplitude away from the equiliibrium position, U, and the kinetic energy, K, and the speed, v, at the same position.
Answer #1

T = 131/68 = 1.93 s

f = 1/T = 0.52 Hz

T = 2pi(m/k)

1.93 = 2pi sqrt(1.09/k)

k = 11.54 N/m

Umax = kx^2/2

Umax = kx^2/2

Umax = 4.58 J

lose in U = gain in KE

max KE = Umax

KEmax = 1/2mv^2

4.58 = 1/2 * 1.09 x v^2

vmax = 2.9 m/s

50.3% of the amplitude away from the equilibrium position is

0.891 x 0.503 = 0.448 m

U = kx^2/2

U = 1.16 J

at that point kinectic energy

KE = Umax - U

KE = 4.58 - 1.16 = 3.42 J

velocity

KE = mv^2/2

3.42 = 1.09 x v^2/2

v = 2.5 m/s